# KANELÉ: Kolmogorov–Arnold Networks for Efficient LUT-based Evaluation

### v1

@@ 1. Introduction
Lookup table (LUT) based neural networks have become a central paradigm for efficient FPGA inference, with designs such as NeuralLUT-Assemble (NeuralLUT_Assemble), TreeLUT (TreeLUT), DWN (DWN), and others (PolyLUT; AmigoLUT) demonstrating dramatic gains in area, latency, and power efficiency. These approaches highlight the advantages of rethinking neural networks around LUT primitives, though they remain largely confined to supervised learning and task-specific architectures.
In this work, we demonstrate that Kolmogorov–Arnold Networks (KANs) offer a principled foundation for LUT-based design. Inspired by the Kolmogorov–Arnold representation theorem, KANs replace the fixed activations of Multilayer Perceptrons (MLPs) with learnable edge functions and the matrix multiplication in MLPs with summation at nodes (Fig. 1). This activation-centric formulation aligns naturally with LUTs: each learnable spline defined on a fixed domain can be quantized, pruned, and directly mapped to LUTs. In the literature, although KANs have been shown to outperform MLPs in settings such as PDE solving and scientific computing (KAN; KAN2), their practical deployment has been hindered by slow inference and costly hardware realizations (Tran_KAN_bad; KAN_BAD_2). The only prior FPGA implementation concluded KANs were impractical, due to expensive spline evaluations and high resource usage (Tran_KAN_bad).
@@
    FPGA-tailored KAN Architecture: We present KANELÉ, named after the French pastry known for its compact form and rich structure (canele). At its core, the framework co-optimizes quantization, pruning, and mapping of KAN functions onto learned LUTs and additions, thereby minimizing memory and logic overhead. From a KAN research perspective, KANELÉ is the first FPGA-tailored formulation, eliminating BRAM/DSP usage, reducing latency by up to $2700\times$, and cutting resource usage by over $4000\times$ compared to prior designs (Tran_KAN_bad).
@@
    Control Systems: we extend KANELÉ beyond supervised learning to continuous control, showing on the HalfCheetah benchmark from OpenAI Gym (Mujoco) that a quantized KAN policy with $\sim 5\times$ fewer parameters than an MLP baseline policy achieves higher rewards, underscoring its suitability for resource-constrained, real-time control systems.
@@ 2.1. Kolmogorov–Arnold Networks
Kolmogorov–Arnold Networks (KANs) replace the fixed activation functions and matrix multiplications of MLPs with learnable spline-based functions on network edges (KAN). This activation-centric formulation improves expressiveness and interpretability, often achieving comparable accuracy with fewer parameters and operations. Since their introduction, KANs have inspired extensive follow-up work, including theoretical analyses (kanormlp2024; expressive2025), architectural extensions (e.g., convolutional (convkan2025; convprinciples2025), temporal (TKAN), and Fourier-based variants (kafn2025; FourierKAN2024)), and applications across scientific modeling and data-driven tasks (Wang_2025; Cruz_2025; li2024ukan; kan4tsf2024).
Despite this rapid progress, recent surveys identify computational efficiency and hardware implementation as key open challenges (survey2024; survey2025; KAN_BAD_2). To date, efficient hardware realization remains largely unexplored, with one early attempt concluding that a direct FPGA implementation incurs prohibitive resource and latency costs compared to MLPs (Tran_KAN_bad). Our work directly challenges this conclusion by demonstrating that the activation-centric design of KANs is, in fact, exceptionally well-suited for hardware acceleration through a LUT-based paradigm.
@@ 2.2. LUT-based Neural Networks
LUT-based neural networks aim to replace arithmetic-heavy MAC operations with precomputed function evaluations stored in LUTs, exploiting the abundant and low-latency logic resources of FPGAs. Pioneering frameworks like LUTNet (LUTNet) and LogicNets (LogicNets) first demonstrated the replacement of arithmetic with direct LUT mappings. Subsequent works generalized this concept to approximate more complex functions (PolyLUT; NeuraLUT) and improved scalability using additive or modular ensembles (PolyLut_Add; AmigoLUT; NeuralLUT_Assemble). This design philosophy has also been used to efficiently implement other machine learning models, such as gradient-boosted decision trees (TreeLUT). Another related approach is the family of Weightless Neural Networks (WNNs), which stores learned patterns directly in LUTs (WISARD; LogicWiSARD; WNN_Edge; WNN_prune; DWN), though often at the cost of representational power.
Conceptually, KANs are close to PolyLUT (PolyLUT), PolyLUT-Add (PolyLut_Add), and DWNs (DWN) but possess distinct structural properties. While PolyLUT tabulates multivariate polynomials, which theoretically allows the native representation of arbitrary products $p(\mathbf{x})=\prod_{i}x_{i}$, this approach suffers from exponential LUT growth relative to input dimension. In contrast, KANs decompose functions into sums of tabulated univariate splines. Although this formulation relies on layer composition to approximate the multiplicative terms inherent to PolyLUT, it yields linear scaling with input dimension and an additive structure that is naturally amenable to pruning. While this formulation doesn’t explicitly represent pure multiplicative terms, in practice, compositions of low-order ($\leq$3) splines approximate such interactions effectively. Moreover, DWN’s full binarization of inputs and LUTs hinders generalization beyond classification, while KANELÉ supports higher-precision arithmetic for tasks such as autoencoding and continuous control. Finally, DWN’s finite-difference differentiability may further constrain optimization flexibility compared to KANELÉ gradient descent.
@@ 3.1. KAN Architecture with Learnable Activation Functions
Before introducing quantization and pruning, we first outline the core architecture. Unlike MLPs, KANs replace fixed nonlinearities with learnable activation functions, each modeled as a linear combination of B-spline basis functions with trainable coefficients. B-splines are piecewise polynomials defined on a grid, providing smoothness, locality, and efficient nonlinear parameterization. More generally, activations can be expressed in other orthogonal bases, such as Fourier series (KAN_GNN; kafn2025; FourierKAN2024).
@@ 3.2. Quantization-Aware Training
For efficient FPGA deployment, we adopt quantization-aware training (QAT) via AMD’s Brevitas library (brevitas). Quantizers are placed at the network input and after each KAN layer, ensuring that training adapts to the required hardware precision.
@@ 3.3. Pruning via Norm-Based Selection
To reduce resource usage, we prune spline connections by evaluating their contribution over the input domain. In contrast to conventional LUT-based neural networks—which rely on sequential LUT indexing, making every LUT entangled with the next and thus nearly impossible to prune without breaking the model—KANELÉ exploits the inherently additive structure of KANs, where each LUT contributes independently to a summation. This independence makes pruning both mathematically natural and directly compatible with FPGA hardware. The original KAN paper emphasizes efficient pruning as a key advantage of KAN’s edge-centric architecture, and we extend this insight to demonstrate a distinct advantage over node-based LUT networks for efficient hardware translation (KAN).
@@ 4. LUT-Based KAN Architecture
This section outlines our end-to-end mapping of trained KANs to synthesizable VHDL RTL and associated pipelining strategies. Our end-to-end toolflow currently supports the basic KAN architecture using B-splines, as in the original paper (KAN). Extensions to other bases or architectures such as convolutions or transformers are feasible. The framework is designed for usability—anyone familiar with training MLPs can readily train and deploy KANELÉ.
@@ 4.1.1. Training
The training process begins by specifying the KAN hyperparameters (see Section 3.4) together with the dataset to be used. The user may freely select the optimizer and learning-rate schedule best suited to their task. In our implementation, we adopt the PyTorch AdamW optimizer (adamW) as the default choice due to its robustness in handling weight decay. The model is then trained with the QAT and pruning mechanism described in Section 3.2 and Section 3.3, respectively, to reduce resource usage and latency, while preserving accuracy. This step produces compact learned activations for KAN that can be efficiently mapped to LUT representations for FPGA deployment.
@@ 4.1.2. KAN to Logical-LUTs Conversion
Following (NeuralLUT_Assemble), we denote lookup tables extracted from the network as Logical-LUTs (L-LUTs), and FPGA fabric resources as Physical-LUTs (P-LUTs). From a trained, pruned, and quantized PyTorch KAN, each surviving connection is translated into an L-LUT. For every active edge, the input state space is enumerated and the KAN layer’s pre-activation response is evaluated and quantized. This produces per-connection truth tables stored as compact JSON files, yielding a deterministic, bit-accurate mapping of the model into integer-valued L-LUTs. The representation preserves quantization and sparsity, enabling efficient FPGA deployment.
@@ 4.2. Pipelining Strategies
As a LUT-based model, KAN inherits known limitations. LUT size scales exponentially with input bitwidth, though only linearly with fan-in (PolyLUT; NeuraLUT). For high-dimensional inputs, preserving KAN’s structure requires long adder chains, reducing throughput and increasing resource usage. Adder trees sustain high clock frequencies by adding pipeline stages, at the cost of extra clock cycles. Thus, image tasks like MNIST require aggressive pruning to remain resource-feasible, with some accuracy loss.
@@ 5.1. Benchmarks
We evaluate KAN across three domains of datasets: (i) widely adopted benchmark datasets from the LUT-based neural network literature, (ii) synthetic and tabular datasets previously used in KAN FPGA benchmarking (Tran_KAN_bad), and (iii) MLPerf Tiny datasets, which provide real-world tasks with more complex modalities and objectives. Together, these domains form a diverse testbed to study KAN across different tasks and levels of complexity. Below we briefly describe each dataset.
@@ 5.1.1. LUT-based Neural Network Benchmarks
  • MNIST: A large-scale handwritten digit recognition dataset containing 60,000 training and 10,000 test images of size $28\times 28$, labeled across 10 classes (digits 0–9) (MNIST).
  • JSC OpenML: A tabular dataset (CERNBox) from the JSC suite, consisting of 16 jet substructure input features and a 5-class jet classification task. It has been widely used in comparisons of LUT-based networks. This version contains 830,000 instances and is known to exhibit easier convergence, possibly due to improved data curation (NeuralLUT_Assemble).
  • JSC CERNBox: Another dataset (OpenML) from the JSC benchmark, involving the same jet tagging task. It comprises 986,806 instances and is generally considered more challenging, as models trained on it tend to achieve lower accuracies due to the increased dataset complexity.
@@ 5.1.2. KAN FPGA Benchmarks
  • Moons: A synthetic two-class dataset (scikit-learn) commonly used for testing nonlinear decision boundaries. Each point lies in one of two interleaving half-moon shapes with added Gaussian noise.
  • Wine: A dataset from the University of California, Irvine (UCI) Machine Learning Repository (wine_dataset), containing 13 physicochemical attributes of wine samples classified into 3 quality categories.
  • Dry Bean: Another UCI dataset (dry_bean) with 16 numerical features representing bean shape and texture, used for classifying 7 different bean varieties .
@@ 5.1.3. MLPerf Tiny Benchmarks
  • ToyADMOS: An audio anomaly detection dataset featuring sound files from both normally functioning and defective toy cars. An autoencoder is trained on sliding windows of the downsampled mel spectrogram (input size 64). For classification, the mean reconstruction loss across all sliding window spectrograms of an audio file is computed, and a fixed threshold is applied to label the sample as an anomaly. Notably, this benchmark presents significantly more complex inputs compared to the LUT-NN and KAN-FPGA benchmarks while also utilizing a non-classification objective (reconstruction loss). (mlperf).
@@
KAN FPGA Benchmarks (Tran_KAN_bad)
@@ 5.3. Comparison with LUT-NN Architectures
We benchmarked KANELÉ against state-of-the-art LUT-based architectures on three datasets: JSC CERNBox, JSC OpenML, and MNIST (Table 3). It should be noted that, as detailed in Table 2, we assume the same input bitwidth compared to prior works, which is not the case for DWN (DWN) that uses a thermometer encoding to assign distinct floating-point thresholds to each feature, leading to potentially large overhead.
@@
NeuraLUT-Assemble (NeuralLUT_Assemble) 75.0 8539 1332 0 0 352 5.7 $4.87\times 10^{4}$
AmigoLUT-NeuraLUT (AmigoLUT) 74.4 42742 4717 0 0 520 9.6 $4.10\times 10^{5}$
PolyLUT-Add (PolyLut_Add) 75.0 36484 1209 0 0 315 16 $5.84\times 10^{5}$
NeuraLUT (NeuraLUT) 75.1 92357 4885 0 0 368 14 $1.29\times 10^{6}$
PolyLUT (PolyLUT) 75.0 246071 12384 0 0 203 25 $6.15\times 10^{6}$
LogicNets (LogicNets) 72.0 37931 810 0 0 427 13 $4.93\times 10^{5}$
@@
NeuraLUT-Assemble (NeuralLUT_Assemble) 76.0 1780 540 0 0 941 2.1 $3.92\times 10^{3}$
TreeLUT (TreeLUT) 75.6 2234 347 0 0 735 2.7 $6.03\times 10^{3}$
DWN (DWN) 76.3 4972 3305 0 0 827 7.3 $9.07\times 10^{4}$
hls4ml (Fahim et al.) (hls4ml_Fahim) 76.2 63251 4394 38 0 200 45 $2.85\times 10^{6}$
@@
NeuraLUT-Assemble (NeuralLUT_Assemble) 97.9 5070 725 0 0 863 2.1 $1.06\times 10^{4}$
TreeLUT (TreeLUT) 96.6 4478 597 0 0 791 2.5 $1.12\times 10^{4}$
DWN (DWN) 97.8 2092 1757 0 0 873 9.2 $1.92\times 10^{4}$
PolyLUT-Add (PolyLut_Add) 96.0 14810 2609 0 0 625 10 $1.48\times 10^{5}$
AmigoLUT-NeuraLUT (AmigoLUT) 95.5 16081 13292 0 0 925 7.6 $1.22\times 10^{5}$
NeuraLUT (NeuraLUT) 96.0 54798 3757 0 0 431 12 $6.58\times 10^{5}$
PolyLUT (PolyLUT) 97.5 75131 4668 0 0 353 17 $1.38\times 10^{6}$
FINN (FINN) 96.0 91131 0 5 200 310 $2.82\times 10^{7}$
hls4ml (Ngadiuba et al.) (hls4ml_Ngadiuba) 95.0 260092 165513 0 345 200 190 $4.94\times 10^{7}$
Table 4. FPGA resource utilization and latency of KAN models on Moons, Wine, and Dry Bean benchmarks as used in (Tran_KAN_bad)
@@
Moons KANELÉ 97 1736 0 0 57 67 5 2.9 $1.7\times 10^{2}$
KAN (Tran et al) (Tran_KAN_bad) 97 - 10 120 8622 17877 128 1280 -
@@
KAN (Tran et al) (Tran_KAN_bad) 97 - 132 950 74741 146843 688 6880 -
@@
KAN (Tran et al) (Tran_KAN_bad) 92 - 781 9111 734544 1677558 1896 18960 -
Table 5. Comparison of FPGA resource utilization, latency, and power consumption for the anomaly detection task on the ToyADMOS time series dataset in the MLPerf Tiny Benchmark, evaluated on the xc7a100t-1csg324 FPGA (mlperf).
@@
hls4ml (MLPerf Tiny v0.7) (Borras_hls4ml) 0.83 22.5 207 61,639 51,429 5,780 144 694 k 45 98.4

### v2

@@ 1. Introduction
Lookup table (LUT) based neural networks have become a central paradigm for efficient FPGA inference, with designs such as NeuralLUT-Assemble (Andronic and Constantinides, 2025), TreeLUT (Khataei and Bazargan, 2025), DWN (Bacellar et al., 2025), and others (Andronic and Constantinides, 2023; Weng et al., 2025; Chang et al., 2024; Sun et al., 2025) demonstrating dramatic gains in area, latency, and power efficiency. These approaches highlight the advantages of rethinking neural networks around LUT primitives, though they remain largely confined to supervised learning and task-specific architectures.
In this work, we demonstrate that Kolmogorov–Arnold Networks (KANs) offer a principled foundation for LUT-based design. Inspired by the Kolmogorov–Arnold representation theorem, KANs replace the fixed activations of Multilayer Perceptrons (MLPs) with learnable edge functions and the matrix multiplication in MLPs with summation at nodes (Fig. 1). This activation-centric formulation aligns naturally with LUTs: each learnable spline defined on a fixed domain can be quantized, pruned, and directly mapped to LUTs. In the literature, although KANs have been shown to outperform MLPs in settings such as PDE solving and scientific computing (Liu et al., 2025, 2024), their practical deployment has been hindered by slow inference and costly hardware realizations (Tran et al., 2024; Hou et al., 2025). The only prior FPGA implementation concluded KANs were impractical, due to expensive spline evaluations and high resource usage (Tran et al., 2024).
@@
    FPGA-tailored KAN Architecture: We present KANELÉ, named after the French pastry known for its compact form and rich structure (Marmiton, n.d.). At its core, the framework co-optimizes quantization, pruning, and mapping of KAN functions onto learned LUTs and additions, thereby minimizing memory and logic overhead. From a KAN research perspective, KANELÉ is the first FPGA-tailored formulation, eliminating BRAM/DSP usage, reducing latency by up to $2700\times$, and cutting resource usage by over $4000\times$ compared to prior designs (Tran et al., 2024).
@@
    Control Systems: we extend KANELÉ beyond supervised learning to continuous control, showing on the HalfCheetah benchmark from OpenAI Gym (Todorov et al., 2012) that a quantized KAN policy with $\sim 5\times$ fewer parameters than an MLP baseline policy achieves higher rewards, underscoring its suitability for resource-constrained, real-time control systems.
@@ 2.1. Kolmogorov–Arnold Networks
Kolmogorov–Arnold Networks (KANs) replace the fixed activation functions and matrix multiplications of MLPs with learnable spline-based functions on network edges (Liu et al., 2025). This activation-centric formulation improves expressiveness and interpretability, often achieving comparable accuracy with fewer parameters and operations. Since their introduction, KANs have inspired extensive follow-up work, including theoretical analyses (Yu et al., 2024; Wang et al., 2025a), architectural extensions (e.g., convolutional (Bodner et al., 2025; Drokin, 2024), temporal (Genet and Inzirillo, 2025), and Fourier-based variants (Zhang et al., 2025; Imran and Ishmam, 2024)), and applications across scientific modeling and data-driven tasks (Wang et al., 2025b; Cruz et al., 2025; Li et al., 2024; Han et al., 2025).
Despite this rapid progress, recent surveys identify computational efficiency and hardware implementation as key open challenges (Ji et al., 2025; Somvanshi et al., 2025; Hou et al., 2025). To date, efficient hardware realization remains largely unexplored, with one early attempt concluding that a direct FPGA implementation incurs prohibitive resource and latency costs compared to MLPs (Tran et al., 2024). Our work directly challenges this conclusion by demonstrating that the activation-centric design of KANs is, in fact, exceptionally well-suited for hardware acceleration through a LUT-based paradigm.
@@ 2.2. LUT-based Neural Networks
LUT-based neural networks aim to replace arithmetic-heavy MAC operations with precomputed function evaluations stored in LUTs, exploiting the abundant and low-latency logic resources of FPGAs. Pioneering frameworks like LUTNet (Wang et al., 2019) and LogicNets (Umuroglu et al., 2020) first demonstrated the replacement of arithmetic with direct LUT mappings. Subsequent works generalized this concept to approximate more complex functions (Andronic and Constantinides, 2023, 2024) and improved scalability using additive or modular ensembles (Lou et al., 2024; Weng et al., 2025; Andronic and Constantinides, 2025). This design philosophy has also been used to efficiently implement other machine learning models, such as gradient-boosted decision trees (Khataei and Bazargan, 2025). Another related approach is the family of Weightless Neural Networks (WNNs), which stores learned patterns directly in LUTs (Aleksander et al., 1984; Miranda et al., 2022; Susskind et al., 2023, 2022; Bacellar et al., 2025), though often at the cost of representational power.
Conceptually, KANs are close to PolyLUT (Andronic and Constantinides, 2023), PolyLUT-Add (Lou et al., 2024), and DWNs (Bacellar et al., 2025) but possess distinct structural properties. While PolyLUT tabulates multivariate polynomials, which theoretically allows the native representation of arbitrary products $p(\mathbf{x})=\prod_{i}x_{i}$, this approach suffers from exponential LUT growth relative to input dimension. In contrast, KANs decompose functions into sums of tabulated univariate splines. Although this formulation relies on layer composition to approximate the multiplicative terms inherent to PolyLUT, it yields linear scaling with input dimension and an additive structure that is naturally amenable to pruning. While this formulation doesn’t explicitly represent pure multiplicative terms, in practice, compositions of low-order ($\leq$3) splines approximate such interactions effectively. Moreover, DWN’s full binarization of inputs and LUTs hinders generalization beyond classification, while KANELÉ supports higher-precision arithmetic for tasks such as autoencoding and continuous control. Finally, DWN’s finite-difference differentiability may further constrain optimization flexibility compared to KANELÉ gradient descent.
@@ 3.1. KAN Architecture with Learnable Activation Functions
Before introducing quantization and pruning, we first outline the core architecture. Unlike MLPs, KANs replace fixed nonlinearities with learnable activation functions, each modeled as a linear combination of B-spline basis functions with trainable coefficients. B-splines are piecewise polynomials defined on a grid, providing smoothness, locality, and efficient nonlinear parameterization. More generally, activations can be expressed in other orthogonal bases, such as Fourier series (Li et al., 2025; Zhang et al., 2025; Imran and Ishmam, 2024).
@@ 3.2. Quantization-Aware Training
For efficient FPGA deployment, we adopt quantization-aware training (QAT) via AMD’s Brevitas library (Franco et al., 2025). Quantizers are placed at the network input and after each KAN layer, ensuring that training adapts to the required hardware precision.
@@ 3.3. Pruning via Norm-Based Selection
To reduce resource usage, we prune spline connections by evaluating their contribution over the input domain. In contrast to conventional LUT-based neural networks—which rely on sequential LUT indexing, making every LUT entangled with the next and thus nearly impossible to prune without breaking the model—KANELÉ exploits the inherently additive structure of KANs, where each LUT contributes independently to a summation. This independence makes pruning both mathematically natural and directly compatible with FPGA hardware. The original KAN paper emphasizes efficient pruning as a key advantage of KAN’s edge-centric architecture, and we extend this insight to demonstrate a distinct advantage over node-based LUT networks for efficient hardware translation (Liu et al., 2025).
@@ 4. LUT-Based KAN Architecture
This section outlines our end-to-end mapping of trained KANs to synthesizable VHDL RTL and associated pipelining strategies. Our end-to-end toolflow currently supports the basic KAN architecture using B-splines, as in the original paper (Liu et al., 2025). Extensions to other bases or architectures such as convolutions or transformers are feasible. The framework is designed for usability—anyone familiar with training MLPs can readily train and deploy KANELÉ.
@@ 4.1.1. Training
The training process begins by specifying the KAN hyperparameters (see Section 3.4) together with the dataset to be used. The user may freely select the optimizer and learning-rate schedule best suited to their task. In our implementation, we adopt the PyTorch AdamW optimizer (Loshchilov and Hutter, 2019) as the default choice due to its robustness in handling weight decay. The model is then trained with the QAT and pruning mechanism described in Section 3.2 and Section 3.3, respectively, to reduce resource usage and latency, while preserving accuracy. This step produces compact learned activations for KAN that can be efficiently mapped to LUT representations for FPGA deployment.
@@ 4.1.2. KAN to Logical-LUTs Conversion
Following (Andronic and Constantinides, 2025), we denote lookup tables extracted from the network as Logical-LUTs (L-LUTs), and FPGA fabric resources as Physical-LUTs (P-LUTs). From a trained, pruned, and quantized PyTorch KAN, each surviving connection is translated into an L-LUT. For every active edge, the input state space is enumerated and the KAN layer’s pre-activation response is evaluated and quantized. This produces per-connection truth tables stored as compact JSON files, yielding a deterministic, bit-accurate mapping of the model into integer-valued L-LUTs. The representation preserves quantization and sparsity, enabling efficient FPGA deployment.
@@ 4.2. Pipelining Strategies
As a LUT-based model, KAN inherits known limitations. LUT size scales exponentially with input bitwidth, though only linearly with fan-in (Andronic and Constantinides, 2023, 2024). For high-dimensional inputs, preserving KAN’s structure requires long adder chains, reducing throughput and increasing resource usage. Adder trees sustain high clock frequencies by adding pipeline stages, at the cost of extra clock cycles. Thus, image tasks like MNIST require aggressive pruning to remain resource-feasible, with some accuracy loss.
@@ 5.1. Benchmarks
We evaluate KAN across three domains of datasets: (i) widely adopted benchmark datasets from the LUT-based neural network literature, (ii) synthetic and tabular datasets previously used in KAN FPGA benchmarking (Tran et al., 2024), and (iii) MLPerf Tiny datasets, which provide real-world tasks with more complex modalities and objectives. Together, these domains form a diverse testbed to study KAN across different tasks and levels of complexity. Below we briefly describe each dataset.
@@ 5.1.1. LUT-based Neural Network Benchmarks
  • MNIST: A large-scale handwritten digit recognition dataset containing 60,000 training and 10,000 test images of size $28\times 28$, labeled across 10 classes (digits 0–9) (Deng, 2012).
  • JSC OpenML: A tabular dataset (CERN Collaboration, 2025) from the JSC suite, consisting of 16 jet substructure input features and a 5-class jet classification task. It has been widely used in comparisons of LUT-based networks. This version contains 830,000 instances and is known to exhibit easier convergence, possibly due to improved data curation (Andronic and Constantinides, 2025).
  • JSC CERNBox: Another dataset (OpenML Contributors and LHC Jets HLF Curators, 2020) from the JSC benchmark, involving the same jet tagging task. It comprises 986,806 instances and is generally considered more challenging, as models trained on it tend to achieve lower accuracies due to the increased dataset complexity.
@@ 5.1.2. KAN FPGA Benchmarks
  • Moons: A synthetic two-class dataset (Pedregosa et al., 2011) commonly used for testing nonlinear decision boundaries. Each point lies in one of two interleaving half-moon shapes with added Gaussian noise.
  • Wine: A dataset from the University of California, Irvine (UCI) Machine Learning Repository (Aeberhard and Forina, 1992), containing 13 physicochemical attributes of wine samples classified into 3 quality categories.
  • Dry Bean: Another UCI dataset (15) with 16 numerical features representing bean shape and texture, used for classifying 7 different bean varieties .
@@ 5.1.3. MLPerf Tiny Benchmarks
  • ToyADMOS: An audio anomaly detection dataset featuring sound files from both normally functioning and defective toy cars. An autoencoder is trained on sliding windows of the downsampled mel spectrogram (input size 64). For classification, the mean reconstruction loss across all sliding window spectrograms of an audio file is computed, and a fixed threshold is applied to label the sample as an anomaly. Notably, this benchmark presents significantly more complex inputs compared to the LUT-NN and KAN-FPGA benchmarks while also utilizing a non-classification objective (reconstruction loss). (Banbury et al., 2021).
@@
KAN FPGA Benchmarks (Tran et al., 2024)
@@ 5.3. Comparison with LUT-NN Architectures
We benchmarked KANELÉ against state-of-the-art LUT-based architectures on three datasets: JSC CERNBox, JSC OpenML, and MNIST (Table 3). It should be noted that, as detailed in Table 2, we assume the same input bitwidth compared to prior works, which is not the case for DWN (Bacellar et al., 2025) that uses a thermometer encoding to assign distinct floating-point thresholds to each feature, leading to potentially large overhead.
@@
NeuraLUT-Assemble (Andronic and Constantinides, 2025) 75.0 8539 1332 0 0 352 5.7 $4.87\times 10^{4}$
AmigoLUT-NeuraLUT (Weng et al., 2025) 74.4 42742 4717 0 0 520 9.6 $4.10\times 10^{5}$
PolyLUT-Add (Lou et al., 2024) 75.0 36484 1209 0 0 315 16 $5.84\times 10^{5}$
NeuraLUT (Andronic and Constantinides, 2024) 75.1 92357 4885 0 0 368 14 $1.29\times 10^{6}$
PolyLUT (Andronic and Constantinides, 2023) 75.0 246071 12384 0 0 203 25 $6.15\times 10^{6}$
LogicNets (Umuroglu et al., 2020) 72.0 37931 810 0 0 427 13 $4.93\times 10^{5}$
@@
NeuraLUT-Assemble (Andronic and Constantinides, 2025) 76.0 1780 540 0 0 941 2.1 $3.92\times 10^{3}$
TreeLUT (Khataei and Bazargan, 2025) 75.6 2234 347 0 0 735 2.7 $6.03\times 10^{3}$
DWN (Bacellar et al., 2025) 76.3 4972 3305 0 0 827 7.3 $3.6\times 10^{4}$
da4ml (Sun et al., 2025) 76.9 12250 1502 0 0 212 18.9 $2.3\times 10^{5}$
hls4ml (Fahim et al.) (Fahim et al., 2021) 76.2 63251 4394 38 0 200 45 $2.85\times 10^{6}$
@@
NeuraLUT-Assemble (Andronic and Constantinides, 2025) 97.9 5070 725 0 0 863 2.1 $1.06\times 10^{4}$
TreeLUT (Khataei and Bazargan, 2025) 96.6 4478 597 0 0 791 2.5 $1.12\times 10^{4}$
DWN (Bacellar et al., 2025) 97.8 2092 1757 0 0 873 9.2 $1.92\times 10^{4}$
PolyLUT-Add (Lou et al., 2024) 96.0 14810 2609 0 0 625 10 $1.48\times 10^{5}$
AmigoLUT-NeuraLUT (Weng et al., 2025) 95.5 16081 13292 0 0 925 7.6 $1.22\times 10^{5}$
NeuraLUT (Andronic and Constantinides, 2024) 96.0 54798 3757 0 0 431 12 $6.58\times 10^{5}$
PolyLUT (Andronic and Constantinides, 2023) 97.5 75131 4668 0 0 353 17 $1.38\times 10^{6}$
FINN (Umuroglu et al., 2017) 96.0 91131 0 5 200 310 $2.82\times 10^{7}$
hls4ml (Ngadiuba et al.) (Ngadiuba et al., 2020) 95.0 260092 165513 0 345 200 190 $4.94\times 10^{7}$
Table 4. FPGA resource utilization and latency of KAN models on Moons, Wine, and Dry Bean benchmarks as used in (Tran et al., 2024)
@@
Moons KANELÉ 97 1736 0 0 57 67 5 2.9 $1.9\times 10^{2}$
KAN (Tran et al) (Tran et al., 2024) 97 - 10 120 8622 17877 128 1280 $2.3\times 10^{7}$
ChebyUnit (Zhonglongyou et al., 2026) 100 - 10 40 12150 9888 13 130 $1.3\times 10^{6}$
@@
KAN (Tran et al) (Tran et al., 2024) 97 - 132 950 74741 146843 688 6880 $1.0\times 10^{9}$
ChebyUnit (Zhonglongyou et al., 2026) 95 - 132 324 22104 30154 13 130 $3.9\times 10^{6}$
@@
KAN (Tran et al) (Tran et al., 2024) 92 - 781 9111 734544 1677558 1896 18960 $3.2\times 10^{10}$
ChebyUnit (Zhonglongyou et al., 2026) 92 - 781 256 25198 27359 13 130 $3.6\times 10^{6}$
Table 5. Comparison of FPGA resource utilization, latency, and power consumption for the anomaly detection task on the ToyADMOS time series dataset in the MLPerf Tiny Benchmark, evaluated on the xc7a100t-1csg324 FPGA (Banbury et al., 2021).
@@
hls4ml (MLPerf Tiny v0.7) (Borras et al., 2022) 0.83 22.5 207 61,639 51,429 5,780 144 694 k 45 98.4